OpenStudy (anonymous):

if the sum of three times a number and negative four is decreased by six then the result is nine less than the original number. Find the original number

OpenStudy (anonymous):

suppose the number is x. the sum of three times a number and negative four : 3x + (-4) = 3x-4 then decreased by six : 3x-4-6 = 3x-10 then the result is nine less than the original number: 3x-10 = x-9 3x-x = 10-9 -> 2x=1 -> x=1/2